Which term of the sequence 3, 12, 48,... is equal to 196 608?




Is it 12th, or, 10th, or 11th, or 9th.

1
Expert's answer
2022-02-17T05:25:04-0500

Solution:

3, 12, 48, ... ,196608

Clearly, the given sequence is an G.P. (Geometric Progression).

The given G.P. has first term a = 3 and the common ratio r = 4.


nth term of a G.P.:

an = a × rn−1


Let the nth term be 196608 (an).

Then,

196608 = 3 × 4n−1

65536 = 4n−1

48 = 4n−1

8 = n − 1

n = 9


Answer: It is 9th term.